1. Calculate the amount of heat required to raise the temperature of 7.25g of liquid water from 22.7oC to 29.2oC.
  2. Calculate the amount of heat required to evaporate 5.20g of liquid water
  3. Calculate the amount of heat to be released when 6.50g of water vapor condenses.
  4. What phenomenon is responsible for the maximum temperature at the boundary of the stratosphere and the atmosphere?
  5. What function does a third body serve in an atmospheric chemical reaction?
  6. Why does the lower boundary of the ionosphere lift at night?
1
Expert's answer
2021-11-02T02:24:26-0400

Is the reaction endothermic or exothermic? Since the temperature of the surrounding increased, that means that heat was generated or given off by the reaction. This makes the reaction EXOTHERMIC. How you know is because the temperature increased


The molar heat of solution is the heat given off when 1 mole of CaCl2 is dissolved in water. The experimental data are for 0.876 g CaCl2, but we will convert to moles later:

heat = q = mC∆T

q = ?

m = mass of water = 95.0 g

C = specific heat of water = 4.184 J/g/deg

∆T = change in temperature = 29.7 - 22.4 = 7.3 degrees

q = (95.0 g)(4.184 J/g/deg)(7.3 deg) = 2902 J

As stated, this is the ∆H for dissolving 0.876 g CaCl2. Now we convert to moles:

0.976 g CaCl2 x 1 mole/110.98 g = 0.00879 moles.


Molar ∆Hrxn = 2902 J / 0.00879 moles = 330,148 J/mole = 330 kJ/mole (to 3 significant figures)